OpenStudy (anonymous):

ln(x)+ln(x-3)is less than or equal to ln4

OpenStudy (anonymous):

ln(x)+ln(x-3)=ln(x*(x-3))=ln(x^2-3) Thus, ln(x)+ln(x-3) is less than ln(4) when x^2<7, and is equal to ln(4) when x^2=7

OpenStudy (anonymous):

that means ln(x^2-3x) less or eq ln4 so x^2 -3x LoE 4

OpenStudy (anonymous):

@Jelliot x^2 - 3x

OpenStudy (anonymous):

sorry bout that... I guess everyone makes mistakes now and again...

OpenStudy (anonymous):

its ok!

OpenStudy (anonymous):